# Towards foundational LiDAR world models with efficient latent flow matching

### v1

@@ 3.3 Improved fine-tuning with representation alignment
To fine-tune the foundational model, we need to feed the data compressor the LiDAR representation from potentially very different domains. We found it important to “align“ the representations of the pre-trained and fine-tuned data prior to use latter to fine-tune the flow matching model. However, in case of subtask 3 (semantic occupancy forecasting), due to the difference in network embedding layer dimensions, it is non-trivial to fine-tune the semantic data based on a VAE pretrained on non-semantic data. Instead, we use the latents of the corresponding dense occupancy $\bm{o}_{d}$ to guide the formulation of the subspace of the semantic VAE.
@@ 4 Experiments
We designed our experiments to investigate the following questions: Q1: Can we improve compression and model performance over existing designs? Q2: Can we develop a LiDAR world model that exhibits substantial superiority to models trained from scratch on three downstream forecasting tasks (high-beams occupancy forecasting, indoor occupancy forecasting, and semantic occupancy forecasting)? Q3: How well does each fine-tuning variant perform and why does fine-tuning the representation work?
@@ 4.1 Experimental design
We use nuScenes (2Hz) as the pre-training data for the flow matching model. We trained 2 different types of foundational models on it, one was trained based on the original lidar sweep ($\bm{o}_{v}$), totaling 27,000 frames, denoted by ${G}_{\theta}^{v}$. The other one was trained based on LiDAR frames after they have been densified ($\bm{o}_{d}$), totaling 19,000 frames, denoted by ${G}_{\theta}^{d}$. The latter training set is a subset of the former part and the number of $\bm{o}_{d}$ equal to $\bm{o}_{s}$.
For the different beam adaptation subtask, we downsampled 11 sequences from the KITTI360 raw dataset liao2022kitti from 10Hz to 2Hz to match the foundational model setting. Also, we collected an indoor navigation dataset using a Clearpath Jackal equipped with an OSO-128 LiDAR sensor (training set with 23,504 frames and validation set with 9,720 frames). Finally, for the semantic occupancy forecasting, we present a 2 fold experiments: in section 4.2, we follow the official splitting tian2023occ3d and train the model from scratch. Then in section and H.1, we only use first half of the training data to pretrain the $G_{\theta}^{d}$ as the foundation model, which then fine-tuning on the other half $\bm{o}_{s}$ ($\bm{o}_{s}$ and $\bm{o}_{d}$ are 1v1 correspondence) to avoid fine-tuning data to be already seen during pretraining stage. From the next section, we use $\bm{o}_{s}^{\prime}$ and $\bm{o}_{d}^{\prime}$ to denote the partial training data. For more details on the model setup, please refer to our appendix.
@@ 4.2 Encoder structure exploration and semantic occupancy forecasting evaluation
For model evaluation, Table 2 compare reconstruction results of our Swin-Transformer VAE with previous methods compress data from 8× to 512×. At 32×, we achieve 99.2% mIoU and 97.9% IoU—far above UniScenes’s 92.1%/87.0% at the same rate. Even at 192×, we maintain 93.9%/85.8%, surpassing OccWorld and DOME by over 11% mIoU, and at an extreme 768× our model still delivers a 9.7% relative gain despite 1.5× higher compression.
Table 1: Comparison of future occupancy forecasting performance on full nuS. validation set. $\dagger$ Methods use future trajectory information. $\star$ DynamicCity generates 16 frames jointly; …
@@
OccWorld zheng2024occworld 25.75 15.14 10.51 17.13 34.63 25.07 20.19 26.63 8.62 1347.09 2.56
RenderWorld yan2024renderworld 28.69 18.89 14.83 20.80 37.74 28.41 24.08 30.08
@@
DynamicCity bian2024dynamiccity 26.18 16.94 34.12 25.82 774.44 19.30
Ours 33.17 21.09 15.64 23.33 40.53 30.37 24.44 31.78 6.29 2.83 389.46 22.22
DOME gu2024dome 29.39 20.98 16.17 22.18 38.84 31.25 26.30 32.13 6.04 5.03 8891.98 0.76
Ours 36.42 27.39 21.66 28.49 43.68 36.89 31.98 37.52 4.55 2.80 389.46 21.43
Table 2: Occupancy reconstruction performance comparison at various compression ratios.
@@
Based on the strong VAE performance, our forecasting model advances the state-of-the-art in semantic occupancy forecasting task and preserves a compact footprint and real time throughput. As shown in Table 1, our model achieves a one-second mIoU of 33.17%, surpassing the previous SOTA model’s 28.69% and OccLLama’s 25.05% by 4.48% and 8.12% respectively and maintaining 21.09% and 15.64% for two and three second prediction which outpaces prior works by at least 2.5% relative performance improvement. Considerable margin is also conveyed in terms of IoU, in average 1.70% higher than previous SOTA model. For future-trajectory-conditioned forecasting, our mIoU surpasses DOME by more than 5.5% across all frames. Critically, these predictive gains from our proposed VAE plus CFM network incur almost no computational overhead. Our semantic occupancy forecasting model runs at 22.22 FPS, requiring only 389.46 GFlops per frame and 30.37 million parameters. Compared to OccWorld, our model achieves around 6% of absolute performance improvement while using only 28.91% of computational cost, less than 50% of the parameters, and almost 10 times faster inference speed. Even for DynamicCity which generates 16 frames altogether, our model uses half of GFlops per frame, 66% of parameters, and 1.1 times higher FPS. In the future-trajectory-conditioned regime, our model sustains 21.4 FPS without increasing in computational cost and parameter count, whereas DOME utilizes 23 times more flops per frame (our model is 4.38% of this amount), 15 times more parameters, 28 times slower FPS, and at least 5.5% less absolute performance. Other metrics: In order to evaluate generative quality of our proposed framework, we report the 3D Fréchet Inception Distance (FID) in Table 1. Moreover, due to the stochastic nature of our proposed CFM model and the limitation of mIoU and IoU in evaluating model’s ability to generate diverse but plausible future predictions, we report the negative log likelihood (NLL) in bits-per-dimension. Please refer to Appendix for more details on parameter counts, metrics computation and performance analysis.
@@ 4.3 LiDAR world model transferability
Using the architecture in Section 4.2 and outdoor-pretrained weights, we evaluate the performance gains from fine-tuning our foundational world model with varying strategies and data scales across individual tasks and present the results in Fig.4. For the high-beam LiDAR adaptation task, all configurations based on the pretrained world model outperform from-scratch VAE and CFM training (blue line) across all frames and data fractions. Pretraining CFM alone yields 17.11% and 6.68% gains for 1s and 3s predictions at 10% data. With both components pretrained and fine-tuned, the gains increase to 14.48% and 26%, respectively. This demonstrates that our Swin Transformer-based VAE effectively captures transferable geometric priors for downstream tasks with varying LiDAR characteristics and the effectiveness of the feature alignment.
@@
For the indoor LiDAR adaptation task, all pretrained model variants outperform from-scratch training when fine-tuning data is limited (less than 25%), with full-parameter tuning of both VAE and CFM achieving the best results. At 10% data usage, full fine-tuning yields at least a 4.47% gain across all prediction horizons. However, as data availability increases (e.g., beyond 25% for 1s prediction), from-scratch training surpasses all fine-tuning methods. This is likely because the indoor dataset is a more distinct task relative to the other tasks, making it easier to learn from scratch without pretrained dynamic priors. We expect this result would no longer hold if substantial indoor data was in the pretraining data. See Appendix for further explanation.
@@
We can also observe similar results in semantic occupancy forecasting task, all VAE/CFM combinations based on the pretrained model outperform training from scratch in mIoU. Notably, at 10% semantic fine-tuning data—which is only 5% of the total used in prior methods—full-parameter fine-tuning yields an 82.6%/80.72%/69.70% relative mIoU gain for 1s/2s/3s prediction, which surpasses OccWorld’s performance zheng2024occworld . After 50% data usage, the performance of from-scratch, exceed that of pretrain CFM-only, and LoRA CFM fine-tuning, while full-parameter VAE+CFM fine-tuning consistently achieves the best performance across all data fractions and prediction horizons.
@@
These results confirm that foundational models can capture transferable LiDAR-based dynamic priors from unlabeled data, supporting downstream tasks requiring semantic interpretation. The improvements are especially notable under low-data regimes, demonstrating reduced dependence on human-labeled samples. Overall, across all tasks, our pretrained model achieves up to 11.17% absolute performance gain, and outperforms training from scratch in 30/36 comparisons points. See Appendix for detailed numerical values corresponding to Figure 4.
@@
However, one question remaining is that why models that have both VAE and CFM modules pretrained (with or without LoRA) consistently possesses higher absolute performance over other settings? In other words, how does fine-turning VAE impact CFM forecasting accuracy?
@@
To further analyse, in Table 3, we compute the similarity of latent spaces under different data scales using CKA kornblith2019similarityneuralnetworkrepresentations and CKNNA huh2024platonic . The similarities are measured between pairwise samples w/w.o semantic information, i.e., $\bm{o}_{d}^{\prime}$ with $\bm{o}_{s}^{\prime}$ obtained from finetuned VAE or VAE trained only on fine-turning data. We observe that the effectiveness of VAE fine-tuning is primarily attributed to aligning the latent space of the new domain with that of the original domain, rather than improving reconstruction accuracy. For example, when using 100% of the fine-turning data $\bm{o}_{s}^{\prime}$, the non-pretrained VAE achieves even higher mIoU/IoU than the pretrained one, yet the final performance shows a clear gap—suggesting that preserving latent structure plays a more crucial role. Although paired data is not available for the first two adaptation scenarios, estimating latent similarity in other domains could further support our hypothesis, details are shown in Appendix.
@@ 4.4 Ablation study
Method U-Net CFM 3D Conv after Temporal DiT CFG Scale mIOU
@@
As mentioned in Section 3.2, we now present a breakdown of the performance improvements contributed by each individual module. As shown in Table 4, we observe consistent gains from CFM, 3D convolution, and CFG components, while using classifier free guidance give us the most significant improvement in both IoU and mIoU. We also experiment the impact of the choice of NFE on FLOPS, FPS, and accuracy. In taking NFE=10, we get the best forecasting accuracy while maitaining considerable level of efficiency in terms of run time and GFlops for semantic occupancy forecasting task. For experiment of VAE data-efficient and other details, please see the Appendix.
@@ 5 Conclusion
In this work, by designing a model that can be efficiently pretrained on large-scale outdoor LiDAR data, we show it can be effectively fine-tuned on diverse downstream tasks. Our approach consistently outperforms training from scratch, especially in low-data regimes. To address inefficiencies in existing models, we introduce a new VAE stucture for high-ratio LiDAR compression and a conditional flow matching approach for forecasting. These components enable strong reconstruction performance and improve training and inference efficiency without sacrificing forecasting accuracy. In addition, our VAE fine-tuning strategies further boost performance. While results are encouraging, future work is needed to extend generalization to additional environments, incorporate multi-modal inputs, and integrate with planning and control systems. Overall, our findings suggest that scalable and transferable LiDAR world models are feasible and can significantly reduce reliance on annotated data in practical applications.

### v2

@@ 3.3 Improved fine-tuning with representation alignment
The pretrain of world model can now be start with framework proposed in 3.2 & 3.1 and large amounts of unlabeled data $\bm{o_{d}}$ in a self-supervised way to learn the unified prior knowledge for environment dynamics. For non-semantic to semantic transfer specifically, we now already have encoder/decoder pair denoted by $q_{d}(\bm{z}_{d}|\bm{o}_{d})$ / $q_{d}(\hat{\bm{o}_{d}}|\bm{z}_{d})$ and forecasting model $G^{d}_{\theta}$. Although we can train the other data compressor $q_{s}(\bm{z}_{s}|\bm{o}_{s})$ / $q_{s}(\hat{\bm{o}_{s}}|\bm{z}_{s})$ and use the corresponding latent $\bm{z}_{s}$ to fine-tuning pretrained $G^{d}_{\theta}$, we have discovered that the latent spaces of $\bm{z}_{s}$ and $\bm{z}_{d}$ are not aligned, which is affects the performance of fine-tuning.
For subtask 1 and 2 in Fig 1, this suboptimality can be tackled by fine-tuning the data compressor first, while in case of subtask 3 (semantic occupancy forecasting), due to the difference in network embedding layer dimensions, it is non-trivial to fine-tune the semantic data based on a VAE pretrained on non-semantic data. Instead, we use the latents of the corresponding dense occupancy $\bm{o}_{d}$ to guide the formulation of the subspace of the semantic VAE.
@@
Therefore, our full pipeline is 3-fold: we first train the data compressor and flow matching model via the easily accessible unlabeled data. Then we use the pretrained data compressor to guide the structure of latent space for subtask’s data. Finally, with the aligned representation, we fine-turn the pretrained weights to obtain the final results.
@@ 4 Experiments
We designed our experiments to investigate several questions: Q1: Can we improve compression and performance over existing designs? Q2: Can we develop a LiDAR world model that exhibits substantial superiority to models trained from scratch on three downstream forecasting tasks (high beam occupancy forecasting, indoor occupancy forecasting, and semantic occupancy forecasting)? Q3: How well does each fine-tuning variant perform and why?
@@ 4.1 Experimental design
We use nuScenescaesar2020nuscenes (2Hz) as the pretraining data for the flow matching model. We trained 2 different types of foundational models on it, one was trained based on the original LiDAR sweep ($\bm{o}_{v}$), totaling 27,000 frames, denoted by ${G}_{\theta}^{v}$. The other type of model was trained using densified LiDAR frames $\bm{o}_{d}$), totaling 19,000 frames, denoted by ${G}_{\theta}^{d}$. The latter training set is a subset of the former part and the number of $\bm{o}_{d}$ equal to $\bm{o}_{s}$.
For the beam adaptation subtask, we down-sampled 11 sequences from the KITTI360 raw dataset liao2022kitti from 10Hz to 2Hz to match the foundational model setting. Also, we collected an indoor navigation dataset using a Clearpath Jackal equipped with an OSO-128 LiDAR sensor (training set with 23,504 frames and validation set with 9,720 frames). Finally, for the semantic occupancy forecasting, we present twofold experiments: in section 4.2, we follow the official splitting tian2023occ3d and train the model from scratch. Then in Section 4.3 and Appendix H.1, we only use first half of the training data to pretrain the $G_{\theta}^{d}$ as the foundation model, which then fine-tuning on the other half $\bm{o}_{s}$ ($\bm{o}_{s}$ and $\bm{o}_{d}$ are 1v1 correspondence) to avoid the dynamic knowledge of fine-tuning data to be already seen during pretraining stage. From the next section, we use $\bm{o}_{s}^{\prime}$ and $\bm{o}_{d}^{\prime}$ to denote the partial training data. For more details on the model setup, please refer to our appendix.
@@ 4.2 Encoder structure exploration and semantic occupancy forecasting evaluation
For model evaluation, Table 2 compare reconstruction results of our Swin‐Transformer VAE with previous methods compress data from 8× to 512×. At 32×, we achieve 99.2% mIoU and 97.9% IoU—far above UniScenes’s 92.1%/87.0% at the same rate. Even at 192×, we maintain 93.9%/85.8%, surpassing OccWorld and DOME by over 11% mIoU, and at an extreme 768× our model still delivers a 9.7% relative gain despite 1.5× higher compression.
Table 1: Comparison of future occupancy forecasting performance on full nuS. validation set. $\dagger$: Methods use future trajectory information. $\star$: DynamicCity generates 16 frames jointly with 1000 steps DDPM-style sampling, we report 50 steps DDIM FPS for faster eval but still keep the original accurancy. All FPS tested based on 1x RTX4090 without kernel fusion or other CUDA acceleration methods. -: Unreported or unable to be computed due to code unavailability.
@@
OccWorld zheng2024occworld 25.75 15.14 10.51 17.13 34.63 25.07 20.19 26.63 72.39 1347.09 16.97
RenderWorld yan2024renderworld 28.69 18.89 14.83 20.80 37.74 28.41 24.08 30.08 416
@@
DynamicCity bian2024dynamiccity 26.18 16.94 34.12 25.82 45.43 774.44 19.30
Ours 33.17 21.09 15.64 23.33 40.53 30.37 24.44 31.78 6.29 30.37 389.46 22.22
DOME gu2024dome 29.39 20.98 16.17 22.18 38.84 31.25 26.30 32.13 6.04 444.07 8891.98 5.48
Ours 36.42 27.39 21.66 28.49 43.68 36.89 31.98 37.52 4.55 30.37 389.46 21.43
Based on the strong VAE performance, our forecasting model advances the state-of-the-art of semantic occupancy forecasting while maintaining real-time throughput. As shown in Table 1, our model achieves a one-second mIoU of 33.17%, surpassing the previous SOTA model’s 28.69% and OccLLama’s 25.05% by 4.48% and 8.12% respectively and maintaining 21.09% and 15.64% for two and three second prediction which outpaces prior works by at least 2.5% relative performance improvement.
Table 2: Occupancy reconstruction performance at various compression ratios.
@@
For future-trajectory-conditioned forecasting, our mIoU surpasses DOME by more than 5.5% across all frames. Our semantic occupancy forecasting model runs at 22.22 FPS, requiring only 389.46 GFlops per frame and 30.37 million parameters. Compared to DynamicCity which generates 16 frames altogether, our model uses half of GFlops per frame, 66% of parameters, and 1.1 times higher FPS. In the future-trajectory-conditioned regime, our model sustains 21.4 FPS without increasing in computational cost and parameter count, whereas DOME utilizes 23 times more flops per frame (our model is 4.38% of this amount), 15 times more parameters, 4 times slower FPS, and at least 5.5% less absolute performance. Due to the stochastic nature of our proposed CFM model and the limitation of mIoU and IoU in evaluating model’s ability to generate diverse but plausible future predictions, we also report the negative log likelihood (NLL) in bits-per-dimension. Our model also present better performance compare with previous SOTA stochastic models.
In order to further measure sample quality, we computed the 3D Fréchet Inception Distance (FID) and Kernel Inception Distance (KID) in Table 4. Our model establishes a new state of the art on both FID and KID. On average, for our model without future trajectories, it reduces FID to just 28.3 % / 48.5 % and KID to 22.7 % / 44.4 % of the scores reported by the deterministic (OccWorld) and stochastic (DOME) models, respectively. See Appendix A.2.2 for more details of these metrics.
Then, benefiting from the low information loss of the data compressor we proposed, it is in fact highly suitable to be used as feature extractor for 4D inception-based metrics (temporal consistency measurement). After we revise the structure with more temporal (4D) module, we retrain the proposed data compressor to eval the FVD score of generated results. In Table 4, to give context for temporal consistency, we include an approach that randomly shuffles the ground truth future ordering and compare the FVD with the correctly ordered GT. This tells us the FVD when the predictions are individually correct but temporally inconsistent. Our model scores of 7.68 and 7.80 are also significantly lower than previous methods.
Table 3: FVD Score Comparison, 3s (6 frames) videos used as input
Method FVD $\downarrow$ ($\times 10^{-3}$)
OceWorld 18.68
DOME 9.79
Ours (hist. traj.) 7.80
Ours (fut. traj.) 7.68
Reorder GT 12.07
Table 4: FID and KID($\times 10^{-2}$) comparison for our method and other generative-based and deterministic-based models.
Method 1s 2s 3s AVG
FID KID FID KID FID KID FID KID
OccWorld 9.54 11.7 8.56 10.46 7.78 9.80 8.62 10.60
DOME 4.37 4.39 5.09 5.44 5.63 6.53 5.03 5.42
Ours (Hist. traj) 1.67 1.49 2.81 2.85 4.02 4.74 2.83 3.02
Ours (Fut. traj) 1.61 1.47 2.90 3.05 3.91 4.55 2.80 3.02
@@ 4.3 LiDAR world model transferability
First, In Table 5, we compare the pretrained model with previous 4D forecasting (occupancy forecasting) methods yang2024visual ; khurana2023point . Ours exhibits better performance in non-semantic occupancy forecasting, particularly at the 3-second horizon. Then, using the architecture in Section 4.2, we evaluate the performance gains from fine-tuning our world model with varying strategies and data scales across individual tasks and present the results in Figure 4.
For the high-beam LiDAR adaptation task, all configurations based on the pretrained world model outperform from-scratch VAE and CFM training (blue line) across all frames and data fractions. Pretraining CFM alone yields 17.11% and 6.68% gains for 1s and 3s predictions at 10% data. With both components pretrained and fine-tuned, the gains increase to 14.48% and 26%, respectively. This demonstrates that our Swin Transformer-based VAE effectively captures transferable geometric priors for downstream tasks with varying LiDAR characteristics and the effectiveness of the feature alignment. For the indoor LiDAR adaptation task, all pretrained model variants outperform from-scratch training when fine-tuning data is limited (less than 25%), with full-parameter tuning of both VAE and CFM achieving the best results. At 10% data usage, full fine-tuning yields at least a 4.47% gain across all prediction horizons.
Table 5: Pretraining performance on different LiDAR representations, without future trajectory. *: ViDAR only released weights trained on $\frac{1}{8}$ dataset.
[…]
1s 2s 3s Avg
ViDAR (I2L)* 13.11 12.48 11.78 12.46
Occ4D (L2L) 26.96 19.51 16.81 21.09
Ours-$\bm{o}_{v}$ 26.98 21.56 18.26 22.27
Ours-$\bm{o}_{d}$ 39.64 29.35 23.73 30.91
@@
However, as data availability increases (e.g., beyond 25% for 1s prediction), training from scratch eventually surpasses all fine-tuning methods. This is likely because the indoor dataset represents a task sufficiently distinct from our pre-training domain, making it easier to learn from scratch without the influence of pre-trained priors. We hypothesize that this result would no longer hold if substantial indoor data were included in the pre-training set. Similar results are also observed in the semantic occupancy forecasting task, where all VAE/CFM combinations based on the pre-trained model outperform training from scratch in mIoU. Notably, with just 10% of the semantic fine-tuning data—which is only 5% of the total data used in prior methods—full-parameter fine-tuning yields an 82.6%/80.72%/69.70% relative mIoU gain for 1s/2s/3s prediction, surpassing OccWorld’s performance zheng2024occworld . After 50% data usage, the performance of from-scratch training exceeds that of the CFM-only and LoRA CFM fine-tuning methods. Nevertheless, full-parameter VAE+CFM fine-tuning consistently achieves the best performance across all data fractions and prediction horizons.
@@
These results confirm that foundational models can capture transferable, LiDAR-based dynamic priors from unlabeled data to support downstream tasks requiring semantic interpretation. The improvements are especially notable in low-data regimes, demonstrating a reduced dependence on human-labeled samples. Overall, across all tasks, our pre-trained model achieves up to an 11.17% absolute performance gain and outperforms training from scratch in 30 out of 36 comparison points. See the Appendix for detailed numerical values corresponding to Figure 4.
@@
However, a key question remains: why do models that fine-tune both the VAE and CFM modules consistently outperform other settings? In other words, how does fine-tuning the VAE impact the CFM’s forecasting accuracy?
@@
To further analyze, in Table 6, we compute the similarity of latent spaces under different data scales using CKA kornblith2019similarityneuralnetworkrepresentations and CKNNA huh2024platonic . The similarities are measured between pairwise samples w/w.o semantic information, i.e., $\bm{o}_{d}^{\prime}$ with $\bm{o}_{s}^{\prime}$ obtained from fine-tuned VAE or VAE trained only on fine-tuning data. We observe that the effectiveness of VAE fine-tuning is primarily attributed to aligning the latent space of the new domain with that of the original domain, rather than improving reconstruction accuracy. For example, when using 100% of the fine-tuning data $\bm{o}_{s}^{\prime}$, the non-pretrained VAE achieves even higher mIoU/IoU than the pretrained one, yet the final performance shows a clear gap—suggesting that preserving latent structure plays a more crucial role. Although paired data is not available for the first two adaptation scenarios, estimating latent similarity in other domains could further support our hypothesis, details are shown in Appendix.
@@ 4.4 Ablation study
Method U-Net CFM
3D Conv
after Temporal DiT
CFG Scale mIoU
@@
As mentioned in Section 3.2, we now present a breakdown of the performance improvements contributed by each individual module. As shown in Table 7, we observe consistent gains from CFM, 3D convolution, and CFG components, while using classifier-free guidance gives us the most significant improvement in both IoU and mIoU. We also examine the impact of NFE on FLOPS, FPS, and accuracy.
As shown in Figure 5, taking NFE=10, we get the best forecasting accuracy while maintaining a considerable level of efficiency in terms of run time and GFlops for the semantic occupancy forecasting task. For the experiment of VAE data-efficient and other details, please see the Appendix.
@@ 5 Conclusion
In this work, by designing a model that can be efficiently pretrained on large-scale outdoor LiDAR data, we show it can be effectively fine-tuned on diverse downstream tasks. Our approach consistently outperforms training from scratch, especially in low-data regimes. To address inefficiencies in existing models, we introduce a new VAE structure for high-ratio LiDAR compression and a conditional flow matching approach for forecasting. These components enable strong reconstruction performance and improve training and inference efficiency without sacrificing forecasting accuracy. In addition, our VAE fine-tuning strategies further boost performance. While results are encouraging, future work is needed to extend generalization to additional environments, incorporate multi-modal inputs, and integrate with planning and control systems. Overall, our findings suggest that scalable and transferable LiDAR world models are feasible and can significantly reduce reliance on annotated data in practical applications.
Acknowledgments
This research was enabled in part by the Digital Research Alliance of Canada (alliancecan.ca).